mod z1 = mod z2 = mod z3 = mod(1/z1 + 1/z2 + 1/z3) = 1 find value of mod( z1 + z2 + z3)

Grade:

2 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

|z_{1}| = |z_{2}| = |z_{3}| = |\frac{1}{z_{1}}+\frac{1}{z_{2}}+\frac{1}{z_{3}}| = 1
|z_{1}|^{2} = 1
z_{1}.z_{1}^{*} = 1
z_{2}.z_{2}^{*} = 1
z_{3}.z_{3}^{*} = 1
|\frac{1}{z_{1}}+\frac{1}{z_{2}}+\frac{1}{z_{3}}| = 1
|z_{1}^{*}+z_{2}^{*}+z_{3}^{*}| = 1
|(z_{1}+z_{2}+z_{3})^{*}| = 1
|z_{1}+z_{2}+z_{3}| = 1


jagdish singh singh
9 years ago
[img]http://latex.codecogs.com/gif.latex?%5Chspace%7B-16%7D%24Given%20%24%5Cbf%7Bx%3D%5Csqrt%7B2%7D%7D%24%5C%5C%5C%5C%5C%5C[/img]